OpenStudy (anonymous):

If 852.04 g Al(NO3)3 reacted with the 741.93 g Na2CO3, which would be the limiting reagent?

OpenStudy (abhisar):

Hello @math92130 ! Can u write the balanced equation for above reaction ?

OpenStudy (anonymous):

yep. 2 Al(NO3)3 + 3Na2CO3 = Al2(CO3)3(s) + 6NaNO3

OpenStudy (abhisar):

So, 2 moles of Al(NO3)3 reqcts with 3 moles Na2CO3

OpenStudy (abhisar):

so how many grams of Al(NO3)3 reacts with how many grams of Na2CO3 ?

OpenStudy (anonymous):

wait shouldn't the mole ratio and molar masses be found

OpenStudy (abhisar):

426 grams of Al(NO3)3 reacts with 318 grams Na2CO3 So 852.04 gram will require 636 grams. Given mass of Na2CO3 is 741.93 grams. So Na2CO3 is not the limitting reagent

OpenStudy (anonymous):

al(no3)3 is the limiting reagent

OpenStudy (abhisar):

yes

OpenStudy (abhisar):

Getting it ?

OpenStudy (abhisar):

According to the balanced equation, 3 moles of Na2CO3 is required to react with 426 grams of al(no3)3. So for 1 gram na2co3 426/318 gram al(no3)3 is required

OpenStudy (abhisar):

Now for 741 g na2co3 426/318 * 741 = 993 gram al(no3)3 is required. But only 852.04 gram is given. So it's a limitting reagent

OpenStudy (anonymous):

okay thanks